OpenStudy (anonymous):

Four times the lesser of two consecutive integers is 12 less than twice the greater number. Find the integer. Does anyone know how to solve that???

OpenStudy (anonymous):

myininaya does

OpenStudy (zarkon):

I don't see it that way

myininaya (myininaya):

oops

OpenStudy (anonymous):

wait now i'm confused :(

OpenStudy (anonymous):

you must be getting tired. i know i am doesn't anyone go out on a saturday night? no wonder the economy is bad

myininaya (myininaya):

satellite don't be mad i accidently deleted yours i meant to delete mine

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

power corrupts!

OpenStudy (anonymous):

lol

myininaya (myininaya):

write how you read it

OpenStudy (zarkon):

that is one way to win a math war

OpenStudy (anonymous):

first number \[n\] second number \[n+1\] four times the first number \[4n\] two times the second \[2(n+1)\] and it is 12 less so you get \[4n=2(n+1)-12\]

myininaya (myininaya):

oh i'm idiot

OpenStudy (anonymous):

i am going to be a moderator and delete all answers above the level of neophyte

myininaya (myininaya):

lesser of the two integers lol

OpenStudy (anonymous):

except zarkon's of course

OpenStudy (zarkon):

ty

OpenStudy (anonymous):

yw

myininaya (myininaya):

i thought i was looking at my name when i was deleting my eyes played tricks on me

OpenStudy (anonymous):

thank you all for helping me. I really needed to get this done so I will probably be online with more questions :D

OpenStudy (anonymous):

post away. as you can see they provide hours of amusement.

OpenStudy (anonymous):

lol wow now i feel really stupid because i have another question

OpenStudy (anonymous):

that was really fast

OpenStudy (anonymous):

my question is...

OpenStudy (anonymous):

two times the least of three consecutive odd integers exceeds three times the greatest by 15 what are the integers???

OpenStudy (anonymous):

your on. i going to bed

OpenStudy (anonymous):

two times the least of three consecutive odd integers exceeds three times the greatest by 15 what are the integers???

OpenStudy (anonymous):

kk bye bye

myininaya (myininaya):

suppose n is odd the next odd integer is n+2 the next next odd integer of n is n+4 n<n+2<n+4 so n is the smaller or the least of these 3 integers so we have 2 times the least of three consecutives has the following representation: 2n

myininaya (myininaya):

and then we have exceeds three times the greatest by 15

myininaya (myininaya):

what is the greatest number here?

myininaya (myininaya):

n n+2 or n+4 ?

OpenStudy (anonymous):

thank you for helping

myininaya (myininaya):

so you answered n+4 right?

myininaya (myininaya):

because that is correct

myininaya (myininaya):

we have 3 times that

myininaya (myininaya):

2n-15=3(n+4) ^ | don't forget about the exceeds by 15 part

OpenStudy (anonymous):

thanks so much!!!!!!!